OK… the Y-arm may be locked with green light, which was the goal, and this is all good but it's not yet awesome. Awesome would be locked and aligned properly and quiet and optimised. So... in order to assist in increasing the awesome-osity, here are a few stream-of-consciousness thoughts and stuff I've noticed and haven't had time to fix/investigate or have otherwise had pointed out to me that may help...
Firstly, the beam is not aligned down the centre of the cavity. It's pretty good horizontally, but vertically it's too low by about 3/4->1cm on ETMY. The mirrors steering the beam into the cavity have no more vertical range left, so in order to get the beam higher the final two mirrors will have to be adjusted on the bench. Adding another mirror to create a square will give more range AND there will be less light lost due to off 45degree incident angles. When I tried this before I couldn't get the beam to return through the Faraday, but now the cavity is properly aligned this should not be a problem.
A side note on alignment - while setting cameras and viewports and things up, Steve noticed that one of the cables to one of the coils (UL) passes behind the ETMY. One of the biggest problems in getting the beam into the system to begin with was missing this cable. It doesn't fall directly into the beam path if the beam is well aligned to the cavity, but for initial alignment it obscures the beam - this may be a problem later for IR alignment.
Next, the final lambda/2 waveplate is not yet in the beam. This will only become a problem when it comes to beating the beams together at the vertex, but it WILL be a problem. Remember to put it in before trying to extract signals for full LSC cavity locking.
Speaking of components and suchlike things, the equipment for the green work was originally stored in 3 plastic boxes which were stored near the end of the X-arm. These boxes, minus the components now used to set up the Y-end, are now similarly stored near the end of the Y-arm.
Mechanical shutter - one needs to be installed on the Y-end just like the X-end. Wasn't necessary for initial locking, but necessary for remote control of the green light on/off.
Other control… the Universal PDH box isn't hooked up to the computers. Connections and such should be identical to the X-arm set-up, but someone who knows what they're doing should hook things up appropriately.
More control - haven't had a chance to optimise the locking and stability so the locking loop, while it appears to be fairly robust, isn't as quiet as we would like. There appears to be more AM coupling than we initially thought based on the Lightwave AM/PM measurements from before. It took a bit of fiddling with the modulation frequency to find a quiet point where the apparent AM effects don't prevent locking. 279kHz is the best point I've found so far. There is still a DC offset component in the feedback that prevents the gain being turned up - unity gain appears limited to about 1kHz maximum. Not sure whether this is due to an offset in the demod signal or from something in the electronics and haven't had time left to check it out properly yet. Again, be aware this may come back to bite you later.
Follow the bouncing spot - the Y-arm suspensions haven't been optimised for damping. I did a little bit of fiddling, but it definitely needs more work. I've roughly aligned the ETMY oplev since that seems to be the mass that's bouncing about most but a bit of work might not go amiss before trusting it to damp anything.
Think that's about all that springs to mind for now…
Thanks to everyone at the 40m lab for helping at various times and answering daft questions, like "Where do you keep your screwdrivers?" or "If I were a spectrum analyser, where would I be?" - it's been most enjoyable!